15 men take 21 days of 8 hours each to do a piece of work. How many days of 6 hours each would 21 women take, if 3 women do as much work as 2 men ?

Answer by richwmiller(17219)

15*21*8=2520
21 women work as 14 men
2520/14=180
6 hour days
180/6=30 days of 6 hours for 21 women